What is the different types and formulas for Profitability Ratios ? (30 marks)
1
Expert's answer
2021-09-29T11:30:52-0400

Solution:

Profitability ratios are financial measurements used by analysts and investors to evaluate and measure the ability of an organization to generate income relative to revenue, operating costs, balance sheet assets, and shareholders’ equity. They display how well a firm utilizes its assets to produce profit and value to shareholders.

 

The profitability ratios are as follows:


Gross profit margin = "\frac{Gross\; profit}{Sales}\times100"


Operating profit margin = "\frac{Operating\; profit}{Sales}\times100"


Net profit margin = "\frac{Net\; Income}{Sales}\times100"


Return on Assets = "\frac{Return\;On \;Equity}{Assets}\times100"


Return on Equity = "\frac{Net\; Income}{Shareholders \; Equity}"